OpenStudy (okdutchman7):

Please help. WILL MEDAL! Mudog can do 8 jobs in 3 days, and Jimmy can do 5 jobs in 2 days. 39 jobs need to be done. Mudog works 3 days and then Jimmy joins in. How many days will both of them have to work together to complete the 39 jobs? @mathmale @sleepyjess @Directrix @mathstudent55 @zepdrix @rebeccaxhawaii

OpenStudy (okdutchman7):

@surjithayer

OpenStudy (kittiwitti1):

How many jobs can each person do per day?

OpenStudy (okdutchman7):

\[Rate _{Mudog}=\frac{ 8 jobs }{ 3 days }\]

OpenStudy (okdutchman7):

\[Rate _{Jimmy}=\frac{ 5 }{ 2 }\]

OpenStudy (okdutchman7):

Rateboth=31/6

OpenStudy (kittiwitti1):

31/6 ??

OpenStudy (okdutchman7):

\[\frac{ 8 }{ 3 } +\frac{ 5 }{ 2 } = \frac{ 31 }{ 6 }\]

OpenStudy (okdutchman7):

That is both of their rates put together.

OpenStudy (kittiwitti1):

Okay, so now you have a rate that you can apply to the amount of jobs that need to be completed.

OpenStudy (okdutchman7):

What do I do next?

OpenStudy (kittiwitti1):

Use this formula: \(\text{rate }(r)\times x\text{ days}=\text{jobs completed in } x\text{ days}\)

OpenStudy (kittiwitti1):

You already have the rate \(r=\frac{31}{6}\), and the jobs they want completed is \(39\). Now, putting this into the equation, you get\[\frac{31}{6}x=39\]

OpenStudy (okdutchman7):

This doesn't work because "Mudog works 3 days and then Jimmy joins in"

OpenStudy (kittiwitti1):

Okay, then calculate how many jobs Mudog completed in 3 days, subtract that number from 39, and calculate again with the new number of jobs instead.

OpenStudy (okdutchman7):

@mathmale What do you think?

OpenStudy (mjdennis):

Don't forget that Mudog works 3 days BEFORE Jimmy joins in! so take away those jobs from the original total, and then use the _method_ that @kittiwitti1 proposed, you'll just need the new number of jobs remaining.

OpenStudy (mathmale):

I've just given kittiwitti1 a medal for 'best explanation," so I'll hold on sharing what I think beyond that.

OpenStudy (kittiwitti1):

Thank you @mathmale (:

OpenStudy (kittiwitti1):

@okdutchman7 are you confused? I can explain again

OpenStudy (okdutchman7):

I'm a little confused kittiwitt. Can you explain it a different way?

OpenStudy (mathmale):

and / or break up the problem into parts.

OpenStudy (kittiwitti1):

Sure. Okay, so Mudog worked 3 days at his rate of 8 jobs in 3 days. Which basically means he finished 8 jobs. Subtract 8 from 39 and you have 31 jobs left to complete. Now we have to use their daily rate, which is \(R_{Mudog}=\frac{8}{3}\) and \(R_{Jimmy}=\frac{5}{2}\). As you've calculated before, these two rates added together get you a composite rate of \(R_{total}=\frac{31}{6}\).

OpenStudy (kittiwitti1):

good so far?

OpenStudy (okdutchman7):

Yes. I understand.

OpenStudy (okdutchman7):

So the answer is 6 days?

OpenStudy (kittiwitti1):

Yes, the answer is 6 days ^_^

OpenStudy (okdutchman7):

Thank you!

OpenStudy (kittiwitti1):

No problem ^_^ If you need any more help on problems I'll be floating around the Math forums (:\[\text{Happy OpenStudying!}\]